OpenStudy (anonymous):

Find the exact value of the trig function given that sin u=5/13 and cos v=-3/5. Both u and v are in the quadrant II: sin (u+v)

OpenStudy (tamtoan):

use the formula for sin(u+v) ...you have everything given

OpenStudy (anonymous):

So I would just do sin ((5/13) + (-3/5))?

OpenStudy (tamtoan):

no, it's in quadrant 2, and you can draw the circle of where you can locate that angle with the sine and cosine given to you. sine = y coordinate and cosine = x coordinate ..then use the formula sin(u+v) = sin(u)cos(v) + sin(v)cos(u)

OpenStudy (anonymous):

Would you be in radian or degree mode for this?

OpenStudy (tamtoan):

something like this ...for cos(v) = -3/5  figure out sine ..then do the same with the other angle.